OpenStudy (anonymous):

Factor 0.04m2 - n2. (0.2n - m)(0.2n + m) (0.2m - n)(0.2m - n) ***** (0.2m - n)(0.2m + n) I think it is the one with *****************

OpenStudy (anonymous):

help me @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's not correct

OpenStudy (anonymous):

okay how do u to it

jimthompson5910 (jim_thompson5910):

think of 0.04m^2 - n^2 as (0.2m)^2 - n^2 this is a difference of squares

jimthompson5910 (jim_thompson5910):

so you would use the difference of squares factoring formula a^2 - b^2 = (a - b)(a + b)

OpenStudy (anonymous):

so witch one is in my answer

OpenStudy (anonymous):

@Shannon20150 , to factor this: 0.04m^(2)-n^(2) okaii first, since the binomial can be factored using the difference of squares formula. This is because both terms are perfect squares. The difference of squares formula is a^(2)-b^(2)=(a-b)(a+b).

OpenStudy (anonymous):

okay

OpenStudy (anonymous):